Starting from the origin, a body oscillates simple harmonically with a period of 2 s. After what time will its kinetic energy be 75% of the total energy?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
½ mv2=3/4(1/2)(vmax)2
A2 ω2cos2 ωt =(3/4)A2ω2
cosωt=\sqrt{3}/2
soωt=\pi /6
t=1/6 sec
Thanks and Regards
Shaik Aasif
askIITians faculty